OpenStudy (anonymous):

t=5s+15 ; s sorry. that's the question

OpenStudy (anonymous):

that's an equation to a graph, you have 2 unknowns

OpenStudy (anonymous):

what does the question ask?

OpenStudy (anonymous):

solve for the given variable: t=5s+15 ; s

OpenStudy (anonymous):

well there are infinite answers to this one, plug any number in to t then solve for s

OpenStudy (anonymous):

or plug any number in to s and solve for t

OpenStudy (anonymous):

this is an example: CP = R ; P CP/R = R/C P = R/C i don't know how to do the problem with numbers though.

OpenStudy (anonymous):

k...maybe some more background info, is this calculus? alg? what grade are you in?